# On the lower bound of the curvature exponent on step-two Carnot groups

### v2

@@ 1.2. Step-two Carnot groups
[…]
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the Lie bracket on $\mathfrak{g}$. Via the Lie group exponential map, we can regard $(\mathbb{G},\cdot)$ as $(\mathfrak{g},\ast)$ with the group operation $\ast$ on $\mathfrak{g}$ given by
@@
Note that on $(\mathfrak{g},\ast)$ the identity is 00 and the inverse element of $a$ is $-a$. For more details, we refer to [9].
@@
[…]
@@ 1.3. Measure contraction property on step-two Carnot groups
where :=$\mathcal{F}:=\{A\subset\mathbb{G}\text{ bounded, measurable with }0<\mu(A)<\infty\}$. Note that from left-invariance, the outer supremum does not play any role and we can choose any $p$ in the group $\mathbb{G}$, in particular the identity element $e$. For an equivalent definition of the geodesic dimension (at a point), we refer to [1, Definition 5.47]. Similarly, in our setting the point will not play any role. The equivalence is a result of [1, Theorem D] there.
@@ 1.4. Main result
The main target of this work is to show that Conjecture 9 is not true as well. However, it is hard to construct a counterexample directly like the one in [24]. Instead, we illustrate this by showing that the new lower bound $N_{0}$ is not lower semicontinuous w.r.t. the convergence of structure constants of Lie algebra (see Definition 10 and Remark 11 below) while the curvature exponent $N_{\mathrm{CE}}$ is. In the following set =$\mathbb{N}=\{1,2,\ldots\}$ and we use $\overline{\mathbb{N}}$ to denote the set ${}\mathbb{N}\cup\{\infty\}$.
@@
2. for any $a,b\in\mathfrak{g}$, we have $[a,b]_{k}\to[a,b]_{\infty}$ as $k\to\infty$.
@@
$[\mathrm{X}_{i},\mathrm{X}_{j}]_{k}=\sum_{\ell=1}^{n-m}c_{ij,k}^{\ell}\mathrm{Y}_{\ell},\qquad\forall\,k\in\overline{\mathbb{N}}.$
@@
For any fixed $k\in\overline{\mathbb{N}}$, the Lie algebra $[\cdot,\cdot]_{k}$ induces a group structure $\ast_{k}$ on $\mathfrak{g}$ by
@@
[…]
@@ 2. The lower bound $N_{0}$
[…]
@@
Furthermore, since $[\mathfrak{g}_{1},\mathfrak{g}_{1}]=\mathfrak{g}_{2}$, for $u\in\mathfrak{g}_{2}\setminus\{0\}$ we must have $J_{u}\neq 0$.
@@ 3. Proof of Theorem 15
[…]
@@
where $k\in\overline{\mathbb{N}}$ and $\frac{1}{\infty}$ is interpreted as 00. In the following for fixed $k\in\overline{\mathbb{N}}$ we use $N_{0,\mathbb{G}_{k}}$, $N_{\mathrm{CE},\mathbb{G}_{k}}$ and $J_{u,\mathbb{G}_{k}}$ to denote the lower bound given in (11), the curvature exponent and the operator defined by (6) on $\mathbb{G}_{k}=(\mathfrak{g},\ast_{k})$ respectively. Note that when $k=\infty$, it is exactly the step-two Carnot group induced by the star graph $K_{1,3}$ in [24, § 7.3] on which $N_{0,\mathbb{G}_{\infty}}=17$. It follows from the construction and Remark 11 that $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ as $k\to\infty$. Now let $\mathfrak{g}_{2}\ni u=\sum_{\ell=1}^{3}u_{\ell}\mathrm{Y}_{\ell}$, under the basis $\{\mathrm{X}_{0},\mathrm{X}_{1},\mathrm{X}_{2},\mathrm{X}_{3}\}$, we have

### v3

@@ 1.2. Step-two Carnot groups
Recall that a connected and simply connected Lie group $\mathbb{G}=(\mathbb{G},\cdot)$ is a step-two Carnot group if its left-invariant Lie algebra $\mathfrak{g}$ admits a stratification
@@
where $\cdot$$\cdot$] denotes the Lie bracket on $\mathfrak{g}$. Via the Lie group exponential map, we can regard $(\mathbb{G},\cdot)$ as $(\mathfrak{g},\ast)$ with the group operation $\ast$ on $\mathfrak{g}$ given by
@@
Note that on $(\mathfrak{g},\ast)$ the identity is 0 and the inverse element of $a$ is a-a. For more details, we refer to [9].
@@
On a step-two Carnot group $\mathbb{G}$, given an inner product $\langle\cdot,\cdot\rangle$ of $\mathfrak{g}_{1}$, the canonical left-invariant sub-Riemannian structure $(\mathcal{D},g)$ is defined as follows: the horizontal distribution $\mathcal{D}$ (a sub-bundle of the tangent bundle $T\mathbb{G}$ satisfying the bracket generating condition) is generated by $\mathfrak{g}_{1}$ and the metric $g$ on $\mathcal{D}$ is determined by the inner product $\langle\cdot,\cdot\rangle$. To be more precise, let $m:=\dim\mathfrak{g}_{1}$ and assume $\{\mathrm{X}_{1},\ldots,\mathrm{X}_{m}\}$ is an orthonormal basis of $\mathfrak{g}_{1}$ under $\langle\cdot,\cdot\rangle$. Then we have
@@ 1.3. Measure contraction property on step-two Carnot groups
where $\mathcal{F}:=\{A\subset\mathbb{G}\text{ bounded, measurable with }0<\mu(A)<\infty\}$. Note that from left-invariance, the outer supremum does not play any role and we can choose any $p$ in the group $\mathbb{G}$, in particular the identity element $e$. For an equivalent definition of the geodesic dimension (at a point), we refer to [1, Definition 5.47]. Similarly, in our setting the point will not play any role. The equivalence is a result of [1, Theorem D] there.
@@ 1.4. Main result
The main target of this work is to show that Conjecture 9 is not true as well. However, it is hard to construct a counterexample directly like the one in [24]. Instead, we illustrate this by showing that the new lower bound $N_{0}$ is not lower semicontinuous w.r.t. the convergence of structure constants of Lie algebra (see Definition 10 and Remark 11 below) while the curvature exponent $N_{\mathrm{CE}}$ is. In the following set $\mathbb{N}=\{1,2,\ldots\}$ and we use $\overline{\mathbb{N}}$ to denote the set $\mathbb{N}\cup\{\infty\}$.
@@
2. for any $a,b\in\mathfrak{g}$, we have $a,b]_{k}\to[a,b]_{\infty}$ as $k\to\infty$.
@@
$\mathrm{X}_{i},\mathrm{X}_{j}]_{k}=\sum_{\ell=1}^{n-m}c_{ij,k}^{\ell}\mathrm{Y}_{\ell},\qquad\forall\,k\in\overline{\mathbb{N}}.$
@@
For any fixed $k\in\overline{\mathbb{N}}$, the Lie algebra $\cdot,\cdot]_{k}$ induces a group structure $\ast_{k}$ on $\mathfrak{g}$ by
@@
In the following we denote $(\mathfrak{g},\ast_{k})$ simply by $\mathbb{G}_{k}$. Furthermore, given some fixed inner product $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}_{1}$ (independent of $k\in\overline{\mathbb{N}}$), we can define a canonical left-invariant sub-Riemannian structure as before and we use $d_{k}$ to denote the Carnot–Carathéodory distance of such sub-Riemannian structure on $\mathbb{G}_{k}$.
@@ 2. The lower bound $N_{0}$
Fix a step-two Carnot group $\mathbb{G}$. To give the explicit formula for $N_{0}$ given in [24] in the setting of step-two Carnot groups, we need to first extend the given inner product $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}_{1}$ to the whole Lie algebra $\mathfrak{g}$ in such way that $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$ are orthogonal. Although different extension will result in different linear maps and spaces (to be defined below), it turns out that the number $N_{0}$ is the same. Since we only focus on the number $N_{0}$ in this work so we can just fix one such inner product on $\mathfrak{g}$. For the result of this subsection, we refer to [24, § 6.4].
@@
Furthermore, since $\mathfrak{g}_{1},\mathfrak{g}_{1}]=\mathfrak{g}_{2}$, for $u\in\mathfrak{g}_{2}\setminus\{0\}$ we must have $J_{u}\neq 0$.
@@ 3. Proof of Theorem 15
On a vector space $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$ with an inner product $\langle\cdot,\cdot\rangle$ w.r.t. which $\mathfrak{g}_{1}$ is orthogonal to $\mathfrak{g}_{2}$, we choose an orthonormal basis of $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$ respectively by
@@
where $k\in\overline{\mathbb{N}}$ and $\frac{1}{\infty}$ is interpreted as 0. In the following for fixed $k\in\overline{\mathbb{N}}$ we use $N_{0,\mathbb{G}_{k}}$, $N_{\mathrm{CE},\mathbb{G}_{k}}$ and $J_{u,\mathbb{G}_{k}}$ to denote the lower bound given in (11), the curvature exponent and the operator defined by (6) on $\mathbb{G}_{k}=(\mathfrak{g},\ast_{k})$ respectively. Note that when $k=\infty$, it is exactly the step-two Carnot group induced by the star graph $K_{1,3}$ in [24, § 7.3] on which $N_{0,\mathbb{G}_{\infty}}=17$. It follows from the construction and Remark 11 that $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ as $k\to\infty$. Now let $\mathfrak{g}_{2}\ni u=\sum_{\ell=1}^{3}u_{\ell}\mathrm{Y}_{\ell}$, under the basis $\{\mathrm{X}_{0},\mathrm{X}_{1},\mathrm{X}_{2},\mathrm{X}_{3}\}$, we have